Activity at the volcano has been at a slightly lower level today, with a reduction in the number of volcanic earthquakes and rockfalls compared to the previous 24 hours. Visibility has been poor.
Between 16:00 yesterday and 16:00 today there were 17 long-period earthquakes, 47 hybrid earthquakes, one volcano-tectonic earthquake and a total of 46 rockfall signals. The hybrids have been generally smaller than over the previous few days. Rockfalls continued throughout the day. Visual observations of the volcano have been restricted by the weather, but glimpses from MVO south around midday indicated that ash and steam venting continues.
The new dome has not been seen for nearly a fortnight now but the seismicity and remote sensing indications are that the dome is still growing. While the activity currently affects mainly the south-west sector of the volcano, dome material remains banked up on the top of the volcano and a rapid switch in the growth area could trigger large pyroclastic flows down any of the ghauts with little warning.
With the heavy rains, the level of ash in the atmosphere in St Peter's and in St John's, as averaged over the last 24 hours, has remained low.
Dr. Ricky Herd has returned to MVO after a short break.